UNIMOLECULAR PROCESSES IN CH_{2}OH BELOW THE DISSOCIATION BARRIER STUDIED BY O-H STRETCH OVERTONE EXCITATION

\underline{J. WEI}, B. KARPICHEV, H. REISLER, Department of Chemistry, University of Southern California, Los Angeles, CA 90089.

Infrared spectra in the fundamental, the first, and the second OH-stretch overtone transitions (ν_{1}, 2ν_{1}, 3ν_{1}) were obtained by using double resonant ionization via the 3p_{z} Rydberg state. The rotationally well-resolved spectra show that intramolecular vibrational redistribution (IVR) is restricted even in the 3ν_{1} energy region around 10 490 cm^{−1}, which is 960 cm^{−1} above the thermochemical threshold for dissociation to H + CH_{2}O. The spectrum in the third overtone region (4ν_{1}) at 13 600 cm^{−1}, which is 3 950 cm^{−1} higher than the dissociation threshold, was obtained by monitoring atomic hydrogen photofragments. The overtone spectrum is partially rotationally resolved with a line width of 1.1 cm^{−1}. The long dissociation lifetime (> 6 ps) and the similar O-H stretch anharmonicties (obtained from the Birge-Sponer plot) for CH_{2}OH (91 cm^{−1}) and other molecules with much higher O-H dissociation energies indicate that the examined energy region is below the barrier to direct O-H bond fission. The lack of signal from deuterium fragments in CD_{2}OH excitation shows that tunneling through the barrier followed by direct O-H bond fission is the dominant pathway following 4ν_{1} excitation. Theoretical calculations predicted that dissociation via isomerization to the methoxy radical, CH_{3}O, has a barrier lower by 1 000 cm^{−1} - 2 000 cm^{−1} than that for direct O-H bond fission, but this channel appears to be unimportant in the 4ν_{1} region. Ongoing work involves accessing higher energy regions that may exceed the reaction barriers. Research funded by DOE.